6 inches width and 126 Square Feet, what is the length ?
Answer by Alan3354(69443)

6 inches width and 126 Square Feet, what is the length ?
----------
Area = L*W = L*(1/2)
126 = L*(1/2)
L/2 = 126
Now what?